### v2

# Context-free Recognition with Transformers

Abstract
Transformers excel empirically on tasks that process well-formed inputs according to some grammar, such as natural language and code. However, it remains unclear how they can process grammatical syntax. In fact, under standard complexity conjectures, standard transformers cannot recognize context-free languages (CFLs), a canonical formalism to describe syntax, or even regular languages, a subclass of CFLs. Past work has shown that ${{\mathcal{O}}}(\log({{n}}))$ looping layers (w.r.t. input length n{{n}}) allows transformers to recognize regular languages, but the question of context-free recognition with looped-transformers remained open. In this work, we show that looped transformers with ${{\mathcal{O}}}(\log({{n}}))$ looping layers and ${{\mathcal{O}}}({{n}}^{6})$ padding tokens can recognize all CFLs. However, training and inference with ${{\mathcal{O}}}({{n}}^{6})$ padding tokens is potentially impractical. Fortunately, we show that, for natural subclasses such as unambiguous CFLs, the recognition problem on transformers becomes more tractable, requiring ${{\mathcal{O}}}({{n}}^{3})$ padding. We empirically validate our results and show that looping helps on a grammar that provably requires logarithmic depth. Overall, our results shed light on the intricacy of CFL recognition by transformers: While general recognition may require an intractable amount of padding, natural constraints such as unambiguity yield efficient recognition algorithms.
1 Introduction
Transformers are proficient at many natural language (qin2024largelanguagemodelsmeet) and coding (jiang2024surveylargelanguagemodels) tasks, both of which involve processing hierarchical structures. Classically, the ability to process hierarchically nested structures is closely connected to the ability to model context-free languages (CFLs). Analysis of internal representations—syntactic probing—has shown that transformers learn to encode syntactic features relevant for parsing, the task of extracting the syntactic structure of a sentence (hewitt-manning-2019-structural; arps2022probingconstituencystructureneural; zhao-etal-2023-transformers). Moreover, empirical work has found that transformers can leverage parallelism to parse distinct substrings of an input in parallel (allen-zhu2025physics; schulz2025unravelingsyntaxlanguagemodels), suggesting an inherent bias towards learning parallel parsing. However, we lack a precise theory of how CFL recognition can be exactly implemented in model internals, and which types of syntax can transformers provably represent. To this end, we study whether transformers can correctly determine the grammaticality of a sentence according to a context-free grammar.
The problem of determining whether an input is grammatical can be stated as the recognition problem for context-free grammars (CFGs): Given a CFG $\mathcal{G}$}}, can a string $\bm{w}$}} be generated by $\mathcal{G}$}}? Several foundational serial parsing algorithms (earley-parsing; cocke; kasami1965AnER; younger) solve this problem. However, such serial procedures cannot be naturally implemented due to transformers’ highly parallel, fixed-depth structure. Even regular languages, a strict subset of CFLs, cannot be recognized by fixed-depth transformers under the standard complexity conjecture ${\textsc{TC}\textsuperscript{0}}\subsetneq{\textsc{NC}\textsuperscript{1}}$ since regular language recognition is complete for NC1 (barrington_monoids) while fixed-depth transformers fall in TC0 (merrill-etal-2022-saturated; chiang2025transformersuniformtc0). Looping layers help: $\log({{n}})$ looping layers (where n{{n}} is the input length) allow transformers to recognize regular languages (merrill-2025littledepthgoeslong). However, the question of whether logarithmic looping enables CFL recognition remains. In this work, we address it by analyzing the difficulty of recognizing various CFL classes by transformers. We conceptualize the difficulty in terms of extra resources needed: looping layers and appending blank padding tokens (merrill2025exactexpressivepowertransformers).
While general CFL recognition cannot be implemented by fixed-depth transformers under standard complexity conjectures, our first result shows via a direct construction that it can be expressed by looping layers ${{\mathcal{O}}}(\log({{n}}))$ times and with ${{\mathcal{O}}}({{n}}^{6})$ padding tokens. To the best of our knowledge, this constitutes the first proof of general CFL recognition by transformers. We then ask whether simpler classes of CFLs can be recognized by transformers with fewer resources. We find that the answer is affirmative: We show that natural subclasses of CFLs can be recognized by simpler transformers. In particular, we identify unambiguity and linearity as two natural properties that make CFL recognition more tractable. Unambiguous CFLs, characterized by strings having at most one possible parse, allow for recognition with reduced padding but more looping. This aligns with transformers’ struggles to parse ambiguous grammars in practice (khalighinejad-etal-2023-approximating). Furthermore, imposing linearity (where each grammar rule has at most one non-terminal on its right-hand side) reduces the amount of looping and padding required for recognizing unambiguous CFLs. We empirically test when looping helps generalization and find it to increase the performance on a CFL known to require ${{\mathcal{O}}}(\log({{n}}))$ time on parallel models of computation, namely the language of variable-free Boolean formulas (buss-bfvp).
In summary, we leverage theory on parallel recognition of CFLs to show that logarithmically-looped transformers can recognize CFLs, characterizing the padding requirements for different relevant subclasses. The results imply that, in contrast to the depth expected to implement a serial algorithm such as CKY, exponentially less depth suffices to recognize general CFLs. While this comes with increased space (padding) requirements in the general case, we show the space can be reduced for natural CFL subclasses. Our main results are summarized in Tab.˜1.
Language class Padding tokens required Looping layers required
General CFLs ${{\mathcal{O}}}({{n}}^{6})$ ${{\mathcal{O}}}(\log({{n}}))$
Unambiguous CFLs ${{\mathcal{O}}}({{n}}^{3})$ ${{\mathcal{O}}}(\log^{2}({{n}}))$
Unambiguous linear CFLs ${{\mathcal{O}}}({{n}}^{2})$ ${{\mathcal{O}}}(\log({{n}}))$
Table 1: The computational resources required by transformers to recognize different classes of context-free languages (CFLs).
2 Preliminaries
An alphabet $\Sigma$}} is a finite, non-empty set of symbols. A string is a finite sequence of symbols from $\Sigma$}}. For a string $\bm{w}$}}, we will denote by n{{n}} the length of the string $\bm{w}$}}. The Kleene closure $\Sigma$$^{*}$}} of $\Sigma$}} is the set of all strings over $\Sigma$}}, and $\varepsilon$}} denotes the empty string. A formal language $\mathbb{L}$} over $\Sigma$}} is a subset of $\Sigma$$^{*}$}}, and a language class is a set of formal languages.
2.1 Context-free Grammars
Definition 2.1.
A context-free grammar (CFG) $\mathcal{G}$}} is a tuple $\left({{\Sigma}},{{\mathcal{N}}},{{{{\mathrm{S}}}}},{{\mathcal{P}}}\right)$}} where: (1) $\Sigma$}}is an alphabet of terminal symbols (2) $\mathcal{N}$}}is a finite non-empty set of nonterminal symbols with ${{\mathcal{N}}}\cap{{\Sigma}}=\emptyset$ (3) ${{\mathcal{P}}}\subseteq{{\mathcal{N}}}\times{{({{\mathcal{N}}}\cup{{\Sigma}})^{*}}}$is a set of production rules of the form ${{\mathrm{A}}}\rightarrow{{{{\bm{\alpha}}}}}$}} for ${{\mathrm{A}}}\in{{\mathcal{N}}}$ and ${{{{\bm{\alpha}}}}}\in{{({{\Sigma}}\cup{{\mathcal{N}}})^{*}}}$ (4) ${{{{\mathrm{S}}}}}\in{{\mathcal{N}}}$is a designated start non-terminal symbol. As standard, we denote terminal and nonterminal symbols by lowercase and uppercase symbols, respectively.
[…]
Definition 2.2.
The language of a CFG $\mathcal{G}$}} is the set ${\mathbb{L}}({{\mathcal{G}}})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\{{{\bm{w}}}\in{{{{\Sigma}}^{*}}}\mid{{{{\mathrm{S}}}}}{{\overset{*}{\Rightarrow}}}{{\bm{w}}}\}$.
Definition 2.3.
A language $\mathbb{L}$} is context-free if there exists a CFG $\mathcal{G}$}} such that ${\mathbb{L}}({{\mathcal{G}}})={\mathbb{L}}$.
It is common practice to consider CFGs in a normal form, namely:
Definition 2.4.
A CFG $\mathcal{G}$}} is in Chomsky Normal Form (CNF) if any ${{p}}\in{{\mathcal{P}}}$ is either of the form ${{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}$}}, ${{\mathrm{A}}}\rightarrow{{a}}$}} or ${{{{\mathrm{S}}}}}\rightarrow{{\varepsilon}}$}}.
Every CFG can be transformed into an equivalent one in CNF.
2.2 Transformers
We consider the idealization of transformers from merrill-2025littledepthgoeslong; merrill2025exactexpressivepowertransformers. In short, [^1] We refer to App. A for more details on the transformer model. we study average hard attention transformers (AHATs), where the attention normalization function returns a uniform average of the values of tokens that maximize the attention score. In practice, training dynamics lead attention heads in transformers to approximate average-hard attention (merrill-etal-2021-effects). The transformers use multi-pre-norm, where the layer normalization is applied before the residual connection on either the entire hidden state or on distinct subsets thereof (merrill2024expressivepowertransformerschain). We further assume logarithmic-precision arithmetic, where computations are performed with ${{\mathcal{O}}}(\log({{n}}))$ bits for an input of size n{{n}}. Coupling AHATs and log-precision unlocks useful gadgets such as storing string indices, counting symbol occurrences across the string and performing equality checks of values stored in residual streams at separate positions (merrill2024expressivepowertransformerschain). We assume input strings to the transformer are augmented with both a beginning-of-sequence (bos) and end-of-sequence (eos) token. Denote by ${{{\bm{x}}}}_{{{\textsc{eos}}}}^{L}$ the contextual representation of eos at end of the forward pass of the transformer. We apply a linear classifier to ${{{\bm{x}}}}_{{{\textsc{eos}}}}^{L}$ to determine string acceptance.
Looped transformers scale the number of layers with input length (merrill-2025littledepthgoeslong).
Definition 2.5.
Let T be a transformer. We denote by $\langle A,B,C\rangle$ a partition of layers such that $A$ is the initial block of layers, $B$ is the looped block of layers and $C$ is the final block of layers. T is d(n)d({{n}})-looped if upon a forward pass with an input of length n{{n}}, $B$ is repeated ${{\mathcal{O}}}(d({{n}}))$ times.
The amount of computation performed by self attention is definitionally quadratic in the string length. One can dynamically increase this by adding padding space (merrill2025exactexpressivepowertransformers).
Definition 2.6.
Let T be a transformer. T is w(n)w({{n}})-padded if ${{\mathcal{O}}}(w({{n}}))$ padding tokens are appended to the end of the string when computing the contextual representations of a length-n{{n}} input.
Scaling number of layers and padding tokens in transformers is analogous to scaling time and space Boolean circuits (merrill2025exactexpressivepowertransformers), a classical parallel model of computation. Allowing for different looping and padding budgets results in different classes of transformers. We adopt naming conventions of these models from merrill2025exactexpressivepowertransformers. We denote by ${{\textsc{AHAT}}}^{d}_{k}$ the class of languages recognized by averaging hard-attention transformers with ${{\mathcal{O}}}(\log^{d}({{n}}))$ looping, ${{\mathcal{O}}}({{n}}^{k})$ padding and strict causal masking. We further denote with uAHAT average hard-attention transformers with no masking, and with mAHAT transformers that use both masked and unmasked attention heads. Conveniently, AHATs can simulate mAHATs:
Lemma 2.1 (merrill2025exactexpressivepowertransformers Proposition 1.).
${{\textsc{uAHAT}}}^{{d}}_{{k}}\subseteq{{\textsc{mAHAT}}}^{{d}}_{{k}}\subseteq{{\textsc{AHAT}}}^{{d}}_{1+\max({{k}},1)}$ for ${{d}}\geq 1$.
3 Recognizing General CFLs with Transformers
We now describe a parallel algorithm for general CFL recognition, which synthesizes ideas from previous work on algorithms for parallel CFL recognition (ruzzo-treesize; rossmanith-observationsparallelrec; lange-unambiguous). We will then show how to implement this algorithm on AHATs, allowing us to prove the following theorem:
Theorem 3.1.
Given a CFL $\mathbb{L}$}, there exists a transformer with both causally-masked and non-masked attention layers, ${{\mathcal{O}}}(\log({{n}}))$ looping layers and ${{\mathcal{O}}}({{n}}^{6})$ padding tokens that recognizes $\mathbb{L}$}. That is, ${\textsc{CFL}}\subseteq{{\textsc{mAHAT}}}^{1}_{6}\subseteq{{\textsc{AHAT}}}^{1}_{7}$.
Our goal is to recognize a CFL represented by a grammar in CNF (Def.˜2.4) with start symbol $\mathrm{S}$}}}}. For a string $\bm{w}$}} of length n{{n}}, the algorithm determines whether ${{\bm{w}}}\in{\mathbb{L}}({{\mathcal{G}}})$. To do this, it manipulates items—tuples of the form ${{\mathrm{A}}},{{i}},{{j}}]$, where ${{\mathrm{A}}}\in{{\mathcal{N}}}$ and ${{i}},{{j}}\in{{\left[{{n}}\right]}}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{\{1,2,\ldots,{{n}}\}}}$. The item ${{\mathrm{A}}},{{i}},{{j}}]$ is realizable if and only if ${{\mathrm{A}}}{{\overset{*}{\Rightarrow}}}{\texttt{w}}_{{i}}{\texttt{w}}_{{{i}}+1}\ldots{\texttt{w}}_{{j}}$, i.e., if there is a sequence of rules that can be applied to the non-terminal $\mathrm{A}$}} that yields ${\texttt{w}}_{{i}}{\texttt{w}}_{{{i}}+1}\ldots{\texttt{w}}_{{j}}$.
We further define slashed items of the form ${{\mathrm{A}}},{{i}},{{j}}]/[{{\mathrm{B}}},{{k}},{{l}}]$, where ${{i}}\leq{{k}}\leq{{l}}\leq{{j}}$. Intuitively, solving ${{\mathrm{A}}},{{i}},{{j}}]/[{{\mathrm{B}}},{{k}},{{l}}]$ equates to determining whether $\mathrm{A}$}} can derive ${\texttt{w}}_{{i}}\ldots{{\mathrm{B}}}\ldots{\texttt{w}}_{{j}}$ assuming that the non-terminal $\mathrm{B}$}} already derives the substring ${\texttt{w}}_{{k}}\ldots{\texttt{w}}_{{l}}$. More formally, ${{\mathrm{A}}},{{i}},{{j}}]/[{{\mathrm{B}}},{{k}},{{l}}]$ is realizable if and only if ${{\mathrm{A}}}{{\overset{*}{\Rightarrow}}}{\texttt{w}}_{{i}}{\texttt{w}}_{{{i}}+1}\ldots{\texttt{w}}_{{{k}}-1}{{\mathrm{B}}}{\texttt{w}}_{{{l}}+1}\dots\texttt{w}_{{{j}}}$.
Naturally, ${{\bm{w}}}\in{\mathbb{L}}({{\mathcal{G}}})$ if and only if the item ${{{{\mathrm{S}}}}},1,{{n}}]$ is realizable, and determining realizability can be broken down recursively as follows:
Lemma 3.1.
${{\mathrm{X}}},{{i}},{{j}}]$ is realizable if and only if one of the following conditions is met:
  • Base case: j=i{{j}}={{i}} and ${{\mathrm{X}}}\rightarrow{\texttt{w}}_{{i}}$}} is a rule in the grammar for some ${\texttt{w}}_{{i}}$.
  • Recursive case 1: There exist a rule ${{\mathrm{X}}}\rightarrow{{\mathrm{Y}}}{{\mathrm{Z}}}$}} and an index k{{k}} such that ${{\mathrm{Y}}},{{i}},{{k}}-1]$ and ${{\mathrm{Z}}},{{k}},{{j}}]$ are realizable items. There are ${{\mathcal{O}}}(|{{\mathcal{P}}}|{{n}})$ ways to choose a rule and an index for ${{\mathcal{O}}}(|{{\mathcal{N}}}|{{n}}^{2})$ possible input items ${{\mathrm{X}}},{{i}},{{j}}]$.
  • Recursive case 2: There exists a ${{\mathrm{Y}}},{{k}},{{l}}]$ such that ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ and ${{\mathrm{Y}}},{{k}},{{l}}]$ are both realizable. There are ${{\mathcal{O}}}(|{{\mathcal{N}}}|{{n}}^{2})$ possible items of the form ${{\mathrm{Y}}},{{k}},{{l}}]$ for ${{\mathcal{O}}}(|{{\mathcal{N}}}|{{n}}^{2})$ possible input items ${{\mathrm{X}}},{{i}},{{j}}]$.
We can also devise a recursive procedure to evaluate the realizability of a slashed items.
Lemma 3.2.
${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ is realizable if and only if one of the following conditions is met:
  • Base case: k=i{{k}}={{i}}, l=j1{{l}}={{j}}-1 and there is a rule ${{\mathrm{X}}}\rightarrow{{\mathrm{Y}}}{{\mathrm{Z}}}$}} in the grammar such that ${{\mathrm{Z}}}\rightarrow{\texttt{w}}_{{j}}$}}. (or the symmetric case)
  • Recursive case 1: There exist a rule ${{\mathrm{X}}}\rightarrow{{\mathrm{A}}}{{\mathrm{B}}}$}} and an index $p$ such that ${{\mathrm{A}}},{{i}},p-1]/[{{\mathrm{Y}}},{{k}},{{l}}]$ and ${{\mathrm{B}}},p,{{j}}]$ are realizable items (or the symmetric case). There are ${{\mathcal{O}}}(|{{\mathcal{P}}}|{{n}})$ ways to choose a rule and an index for ${{\mathcal{O}}}(|{{\mathcal{N}}}|^{2}{{n}}^{4})$ possible input slashed items ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$.
  • Recursive case 2: There exists a ${{\mathrm{Z}}},p,q]$ such that ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Z}}},p,q]$ and ${{\mathrm{Z}}},p,q]/[{{\mathrm{Y}}},{{k}},{{l}}]$ are both realizable. There are ${{\mathcal{O}}}(|{{\mathcal{N}}}|{{n}}^{2})$ possible items of the form ${{\mathrm{Z}}},p,q]$ for ${{\mathcal{O}}}(|{{\mathcal{N}}}|^{2}{{n}}^{4})$ possible input slashed items ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$.
Proof.
Lem.˜3.1 and Lem.˜3.2 both follow from first principles. In the base case of ${{\mathrm{X}}},{{i}},{{j}}]$, if j=i{{j}}={{i}}, then $\mathrm{X}$}} needs to derive exactly the symbol ${\texttt{w}}_{{i}}$ in one step without producing non-terminals (assuming a CFG with no useless non-terminals) because the grammar is in CNF. Similarly, in the base case of ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$, $\mathrm{X}$}} needs to derive in one step a non-terminal $\mathrm{Y}$}} and a non-terminal $\mathrm{Z}$}} such that $\mathrm{Z}$}} derives in one step a symbol at the boundary of ${\texttt{w}}_{{i}}\ldots{\texttt{w}}_{{j}}$. In the recursive case, a (slashed) item exists if and only if it can be described by a parse tree. Such a parse tree can be broken down into recursive subproblems by selecting a split node, which induces two sub-items to solve. We distinguish two equivalent cases: selecting the root as the split (which requires replacing the root $\mathrm{X}$}} with non-terminals it can produce in one-step) [^2] This is the same recursive rule as allen-zhu2025physics’s recursive formula., or a non-root node as the split. ∎
Parallel algorithms for CFL recognition.
Lemmata˜3.1 and 3.2 state that an item is realizable if it can be decomposed into realizable subproblems. Rather than enumerating all the possible decompositions sequentially, we will leverage parallelism to simultaneously compute the realizability of all the induced subproblems. The term guessing has been coined (chandra-alternation) to denote the ability of a parallel model of computation to attend to a valid computation path given an unbounded set of possible computations via parallel branching. By analogy, we can guess which of the correct decompositions of an item is correct by leveraging parallelism, and then recursively verify the induced subproblems in parallel. This suggests natural parallel algorithms for checking realizability, which we present in Algs.˜1 and 2.
Algorithm 1 Determining if the item ${{\mathrm{X}}},{{i}},{{j}}]$ is realizable.
[…]
Algorithm 2 Determining if the item ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ is realizable.
[…]
Intuitively, the recursive function Solve defined in Algs.˜1 and 2 computes the realizability of items.
Theorem 3.2 (Correctness).
Given a CFG $\mathcal{G}$}} in CNF and ${{\bm{w}}}\in{{{{\Sigma}}^{*}}}$ of length $n$, $\textsc{Solve}([{{{{\mathrm{S}}}}},1,{{n}}])={\mathrm{T}}$ if and only if ${{\bm{w}}}\in{\mathbb{L}}({{\mathcal{G}}})$.
Proof.
By definition, ${{\bm{w}}}\in{\mathbb{L}}({{\mathcal{G}}})$ if and only if ${{{{\mathrm{S}}}}},1,{{n}}]$ is realizable. By Lemmata˜3.1 and 3.2, the item ${{{{\mathrm{S}}}}},1,{{n}}]$ is realizable if and only if there exists a decomposition of ${{{{\mathrm{S}}}}},1,{{n}}]$ that respects Lemmata˜3.1 and 3.2. Solve recursively computes all such decompositions, guaranteeing that we will encounter a valid decomposition if it exists. ∎
We now analyze the resources required to compute $\textsc{Solve}[{{{{\mathrm{S}}}}},1,{{n}}]$, which is equivalent to testing membership of the input string $\bm{w}$}} in the given grammar $\mathcal{G}$}}. The recursive procedure induced by Solve is based on a balanced decomposition of problems into subproblems of roughly equal size, which intuitively leads to a $\log({{n}})$-time procedure. Formally, we have the following well-known theorem for decomposing trees:
Theorem 3.3 (jordan1869).
Given a tree with n{{n}} nodes, there exists a node whose removal partitions the tree into two trees with each at most n/2{{n}}/2 nodes.
We rely on Thm.˜3.3 to prove that Alg.˜1 runs in a logarithmic number of recursive steps:
Theorem 3.4 (Parallel Runtime).
We can compute $\textsc{Solve}([{{{{\mathrm{S}}}}},1,{{n}}])$ in $\log({{n}})+{{\mathcal{O}}}(1)$ recursive steps $\forall{{\bm{w}}}\in{{{{\Sigma}}^{*}}}$ with $|{{\bm{w}}}|={{n}}$.
Proof.
By Thm.˜3.3, for any realizable item, there exists a balanced decomposition of the corresponding parse tree into two trees of roughly equal size which can be represented by two items (the split is at the root) or a slashed item and an item (the split is not at the root). Assuming we can process all possible tree decompositions in parallel, we will necessarily guess the balanced one where subtrees have at most 2n/2+12{{n}}/2+1 nodes (a full binary tree with n{{n}} leaves does not have more than 2n{{n}} nodes). After i{{i}} recursive steps, the current subtrees have at most $\frac{{{n}}}{2^{{{i}}-1}}+{{\mathcal{O}}}(1)$ nodes. Therefore, we will solve all items after at most $\log({{n}})+{{\mathcal{O}}}(1)$ steps. ∎
Space complexity. The bottleneck resides in solving an item ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$, which occupies ${{\mathcal{O}}}({{n}}^{4})$ space, and guessing an item ${{\mathrm{Z}}},p,q]$ that could decompose this problem, which itself occupies ${{\mathcal{O}}}({{n}}^{2})$ space, leading to a total space complexity of ${{\mathcal{O}}}({{n}}^{6})$. Combining both insights on time- and space-complexity, we can then prove the following theorem: See 3.1
Proof intuition.
The construction implements Algs.˜1 and 2 on a transformer. Intuitively, each item and possible decomposition is associated with a padding token. There are ${{\mathcal{O}}}({{n}}^{6})$ ways to enumerate items and a possible decomposition. We assume a three-value logic system, where each item is associated with a value in $\{0,1,\bot\}$}} to denote that the item is unrealizable (0), realizable (1) or not known yet to be realizable ($\bot$ [^3] We write $\bot$ for ease of notation. Concretely, $\bot$ can be encoded as any integer that is neither 0 nor 1.). Each padding token allocates space for this value. Intuitively, we will develop a construction such that padding tokens compute the information of whether their associated item is realizable w.r.t. the given decomposition. Initially, all padding tokens store $\bot$. In the initial block of layers, padding tokens associated with a base case item of the form ${{\mathrm{A}}},{{i}},{{i}}]$ can attend to symbol representations via an equality-check to verify whether the base case is valid, i.e., ${{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}$. In the inductive step, padding tokens attend to the padding tokens associated with the decomposition via an equality-check. A feedforward network then either adds $1$ to the residual stream if both sub-items are realizable, 0 if any of them is non-realizable, or $\bot$ if realizability can not be determined at the current iteration. It takes $\log({{n}})$ looping layers to populate the values of all items in their respective padding tokens due to Thm.˜3.3. Finally, we can check whether there exists a padding token associated with ${{{{\mathrm{S}}}}},1,{{n}}]$ that holds the value $1$. Applying Lem.˜2.1 yields inclusion in ${{\textsc{AHAT}}}^{1}_{7}$. The detailed proof is in §˜B.1. ∎
4 Unambiguity Reduces Padding Requirements for Recognition
§˜3 shows that $\log({{n}})$-depth mAHATs with ${{\mathcal{O}}}({{n}}^{6})$ padding can recognize all CFLs. Intuitively, the role of padding in our construction is to handle ambiguity in an arbitrary CFL by storing all the ways in which we can decompose an item. Guessing how to decompose an arbitrary item seemingly requires a substantial amount of space. Therefore, one might conjecture that constraining a grammar to be less ambiguous could potentially reduce the space requirements for recognition. Accordingly, we next study unambiguous CFLs, where there is at most one possible derivation (i.e., parse tree) for any input string. We show will recognizing unambiguous CFLs requires less padding via the following theorem.
Theorem 4.1.
Let UCFL be the class of unambiguous CFLs. Then ${\textsc{UCFL}}\subseteq{{\textsc{mAHAT}}}^{2}_{3}\subseteq{{\textsc{AHAT}}}^{2}_{4}$.
Unambiguity is a natural CFL feature of general interest. Transformers struggle to parse ambiguous grammars (khalighinejad-etal-2023-approximating) and struggle to process syntactically ambiguous natural language sentences (liu2023wereafraidlanguagemodels). Moreover, modern parsers for programming languages such as LR parsers rely on deterministic (therefore unambiguous) CFLs to process inputs in linear time.
This section first introduces an unambiguous CFG recognition algorithm with a tractable space complexity in $\log^{2}({{n}})$-time. We then translate this algorithm into AHATs with a tractable amount of padding.
4.1 A Path System Framework for Unambiguous CFL Recognition
We formalize recognition of unambiguous CFLs as a path system problem (chytil-unambiguous). A path system consists of initial nodes that are associated with either the value $\mathrm{T}$} or $\mathrm{F}$}, and a relation $\mathcal{R}$}}} that specifies how to connect the nodes. By associating base case items of the form ${{\mathrm{A}}},{{i}},{{i}}]$ to initial nodes, general items of the form ${{\mathrm{A}}},{{i}},{{j}}]$ to arbitrary nodes, and connecting nodes depending on the rules of the given grammar, we can compute the realizability of an item by finding a path between its associated nodes and a base node. We now present chytil-unambiguous’s path system framework for recognizing unambiguous CNF CFGs and express it in AHATs.
We denote by $\mathcal{V}$} a set of nodes, each associated with a tuple ${{\mathrm{A}}},{{i}},{{j}}]$. We denote by ${\mathcal{T}}\subseteq{\mathcal{V}}$ the initial set of nodes of the form ${{\mathrm{A}}},{{i}},{{i}}]$ such that ${{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}$. ${\mathcal{R}({{{x}}},{{{y}}},{{{z}}})}\colon{\mathcal{V}}^{3}\to{{\{0,1\}}}$ is a function that describes how to relate the nodes, where ${\mathcal{R}({{{x}}},{{{y}}},{{{z}}})}={\mathrm{T}}$ if and only if z{{{z}}} is associated with some tuple ${{\mathrm{A}}},{{i}},{{j}}]$, x{{{x}}} is associated with some tuple ${{\mathrm{B}}},{{i}},{{k}}]$, and y{{{y}}} is associated with some tuple ${{\mathrm{C}}},{{k}},{{j}}]$ such that ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$. We denote by ${\mathcal{C}}({{\bm{w}}})\subseteq{\mathcal{V}}$ the smallest set containing $\mathcal{T}$} such that if ${{{x}}},{{{y}}}\in{\mathcal{C}}({{\bm{w}}})$ and ${\mathcal{R}({{{x}}},{{{y}}},{{{z}}})}={\mathrm{T}}$ then ${{{z}}}\in{\mathcal{C}}({{\bm{w}}})$, i.e., ${\mathcal{C}}({{\bm{w}}})$ is the closure of $\mathcal{T}$} with respect to $\mathcal{R}$}}}. Equivalently, ${\mathcal{C}}({{\bm{w}}})$ is defined as the set of realizable elements, and the recognition problem is thus equivalent to determining whether the node associated with ${{{{\mathrm{S}}}}},1,{{n}}]$ is in the set ${\mathcal{C}}({{\bm{w}}})$.
Let us now describe how to compute ${\mathcal{C}}({{\bm{w}}})$. Let ${{{\mathcal{X}}}}\subseteq{\mathcal{V}}$ be a set of marked nodes. A dependency graph with respect to $\mathcal{X}$}}}, denoted ${\textsc{DG}}({{{\mathcal{X}}}})$, is the directed graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ where:
${\mathcal{E}}$ $=\{({{{z}}},{{{x}}})\mid{{{z}}}\notin{{{\mathcal{X}}}},{\mathcal{R}({{{x}}},{{{y}}},{{{z}}})}={\mathrm{T}}$ (1)
$\text{ or }{\mathcal{R}({{{y}}},{{{x}}},{{{z}}})}={\mathrm{T}}\text{ for some }{{{y}}}\in{{{\mathcal{X}}}}\}$ (2)
Intuitively, assuming ${{{\mathcal{X}}}}\subseteq{\mathcal{C}}({{\bm{w}}})$, the edge (z,x)({{{z}}},{{{x}}}) can be interpreted as follows: ${{{x}}}\in{\mathcal{C}}({{\bm{w}}})$ implies that ${{{z}}}\in{\mathcal{C}}({{\bm{w}}})$. Precisely, (z,x)({{{z}}},{{{x}}}) being an edge signals that there is some node y{{{y}}} associated with a realizable item such that ${\mathcal{R}({{{x}}},{{{y}}},{{{z}}})}={\mathrm{T}}$ or ${\mathcal{R}({{{y}}},{{{x}}},{{{z}}})}={\mathrm{T}}$. Therefore, if x{{{x}}} is also associated with a realizable item (i.e, is in the closure ${\mathcal{C}}({{\bm{w}}})$), then z{{{z}}} is a realizable item. The algorithm iteratively expands the known set of nodes to be associated with realizable items by computing the set of nodes that have a directed path to a marked node. We denote by Reach($\mathcal{D}$}}}) the nodes of the dependency graph $\mathcal{G}$} that have a directed path to a marked node in $\mathcal{D}$}}}. chytil-unambiguous’s procedure to compute ${\mathcal{C}}({{\bm{w}}})$ is described in Alg.˜3.
$\textsc{Compute closure}({{\bm{w}}},{{\mathcal{G}}})$${\mathcal{V}}\leftarrow{{\{[{{\mathrm{A}}},{{i}},{{j}}]\}}}initialize$${\mathcal{T}}\leftarrow{{\{[{{\mathrm{A}}},{{i}},{{i}}]\mid{{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}\}}}$${{{\mathcal{X}}}}\leftarrow{\mathcal{T}}$_ in range $\log({{n}})$${{{\mathcal{D}}}}\leftarrow{\textsc{DG}}({{{\mathcal{X}}}})$${{{\mathcal{X}}}}\leftarrow\textsc{Reach}({{{\mathcal{D}}}})$$\mathcal{X}$}}}
[…]
Fact 4.1 (chytil-unambiguous).
Let $\mathbb{L}$} be an unambiguous CFL, let $\mathcal{G}$} be a corresponding dependency graph as defined in Eq.˜1. Then, for any pair of nodes in $\mathcal{G}$}, there is at most one directed path from one node to the other.
Therefore, for each node $v$, the subgraph induced by nodes reachable from $v$ becomes a tree rooted at $v$. Reachability queries on a tree reduce to evaluating the corresponding Boolean formula, where leaf nodes are assigned $\mathrm{T}$} if they correspond to realizable items and non-leaf nodes are assigned the $\lor$ operator. We rely on the following lemma to perform this procedure on transformers:
Lemma 4.1.
Let $\psi$ be a variable-free Boolean formula. Consider the binary expression tree of $\psi$, denoted by ${\mathcal{G}}_{\psi}$. Assume all subformulas of $\psi$ are represented in a transformer’s residual stream as follows. For each leaf of ${\mathcal{G}}_{\psi}$, there is a token that encodes its value ($\mathrm{T}$} or $\mathrm{F}$}). For each function node of ${\mathcal{G}}_{\psi}$, there is a token that encodes its type ($\land$ or $\lor$) and pointers to its input arguments. Then, there exists a ${{\mathcal{O}}}(\log({{n}}))$-looped transformer that adds the values of each subformulas in their associated tokens’ residual stream.
Proof intuition.
[…]
We denote by BFVP the set of of variable-free Boolean formulas that evaluate to $\mathrm{T}$}. Importantly, it is known that BFVP is a NC1-complete, unambiguous CFL (buss-bfvp). In other words, BFVP is known to require ${{\mathcal{O}}}(\log({{n}}))$-time on classical models of parallel computation. As a by-product of Lem.˜4.1, we obtain as a free result that $\log$-depth transformers with no padding can recognize BFVP.
Corollary 4.1.
${\textsc{BFVP}}\in{{\textsc{AHAT}}}^{1}_{0}$.
Proof.
See §˜B.2. ∎
We can now show how to simulate Alg.˜3’s procedure on transformers for unambiguous CFLs with ${{\mathcal{O}}}(\log({{n}})^{2})$ looping layers and ${{\mathcal{O}}}({{n}}^{3})$ padding tokens.
See 4.1
Proof intuition.
We implement Alg.˜3 on mAHATs. Each item ${{\mathrm{A}}},{{i}},{{j}}]$ (of which there are ${{\mathcal{O}}}({{n}}^{2})$) is assigned a padding token. For each item ${{\mathrm{A}}},{{i}},{{j}}]$, there are ${{\mathcal{O}}}({{n}})$ ways to decompose it using a split index ${{k}}\in{{\left[{{n}}\right]}}$. For every potential edge between nodes associated with ${{\mathrm{A}}},{{i}},{{j}}]$ and some ${{\mathrm{B}}},{{i}},{{k}}]$ (or ${{\mathrm{B}}},{{k}},{{j}}]$), we assign a padding token. As in Thm.˜3.1, we assume a three-valued logic system where padding tokens for nodes are at any step assigned an element in $\{0,1,\bot\}$}}, denoting non-realizability (0), realizability (1) or not yet known to be realizable ($\bot$). Initially, all padding tokens store $\bot$.
Initially, padding tokens for nodes can check whether they are associated with base case items of the form ${{\mathrm{A}}},{{i}},{{i}}]$. These padding tokens can add to their residual stream $1$ (item is realizable) or 0 (item is non-realizable) depending on if ${{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}$.
In the iterative case, each padding token for an edge associated with items ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$ can first check whether there exists a rule ${{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}$}} and if so, add to the residual stream ${{\mathrm{C}}},{{k}}+1,{{j}}]$. Crucially, there are finitely such items (proportional to $|{{\mathcal{N}}}|$ as the splitting index k{{k}} is fixed). Padding token for edges can attend to padding tokens associated with ${{\mathrm{C}}},{{k}}+1,{{j}}]$ and check whether any of them stores 1, denoting realizability. In that case, the padding token associated with ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$ signals that the edge (${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$) is now in the graph (following how we define edges in Eq.˜1). Padding tokens for nodes associated with items ${{\mathrm{A}}},{{i}},{{j}}]$ can therefore attend to padding tokens for edges associated with ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$, which yields the dependency graph.
Crucially, due to unambiguity, for each node $v$, the subgraph induced by nodes reachable from $v$ becomes a tree rooted at $v$ (Fact˜4.1). We then show how to binarize this tree by extending it with intermediary nodes and edges. Reachability queries on a binary tree can be reduced to the evaluation of the induced Boolean formula (chytil-unambiguous). We invoke Lem.˜4.1 to evaluate Boolean formulas in $\log({{n}})$ steps. The detailed proof is in §˜B.2. ∎
Table 2: Mean accuracy ($\pm$ standard deviation) by language and transformer type across seeds.
Test accuracy on in-distribution strings Test accuracy on out-of-distribution strings
Language Fixed-depth $\log({{n}})$ looping Fixed-depth $\log({{n}})$ looping
BFVP $0.97_{\pm 0.01}$ $0.98_{\pm 0.00}$ $0.88_{\pm 0.01}$ $0.91_{\pm 0.01}$
BFVP (postfix) $0.95_{\pm 0.01}$ $0.98_{\pm 0.00}$ $0.87_{\pm 0.01}$ $0.91_{\pm 0.01}$
Palindrome $0.94_{\pm 0.01}$ $0.93_{\pm 0.01}$ $0.79_{\pm 0.03}$ $0.72_{\pm 0.03}$
Marked palindrome $0.97_{\pm 0.01}$ $0.98_{\pm 0.01}$ $0.59_{\pm 0.19}$ $0.66_{\pm 0.18}$
$\mathrm{D}\!\left(1\right)$}} $0.98_{\pm 0.00}$ $0.98_{\pm 0.00}$ $0.94_{\pm 0.02}$ $0.93_{\pm 0.01}$
$\mathrm{D}\!\left(2\right)$}} $8_{\pm 0.02} 0.99±0$$9_{\pm 0.00}$ $0.83_{\pm 0.08}$ $0.90_{\pm 0.08}$
4.2 Unambiguous Linear CFLs Require Less Time and Space
Finally, we show how linearity further reduces the resources needed by transformers to recognize unambiguous CFLs. A linear CFL is one recognized by a CFG where each rule is the form ${{\mathrm{A}}}\rightarrow{{a}}{{\mathrm{B}}}$}}, ${{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{a}}$}}, or ${{\mathrm{A}}}\rightarrow{{a}}$}}. While restricted, linear CFLs capture a wide range of features of context-freeness. For example, balanced counting can be modeled by the linear CFL ${\mathbb{L}}=\{{{a}}^{{n}}{{b}}^{{n}}\mid{{n}}\geq 0\}$, and symmetry can be modeled by the linear CFL ${\mathbb{L}}=\{{{\bm{w}}}{{\bm{w}}}^{R}\mid{{\bm{w}}}\in{{{{\Sigma}}^{*}}}\}$.
We consider unambiguous linear [^4] There is a subtlety here: A CFL can be induced by both a non-linear unambiguous grammar and by a different linear, ambiguous grammar. Here we consider grammars that are simultaneously linear and unambiguous. CFLs (ULCFLs) and show they can be recognized by $\log$-depth transformers with quadratic padding.
Theorem 4.2.
${\textsc{ULCFL}}\subseteq{{\textsc{mAHAT}}}^{1}_{2}\subseteq{{\textsc{AHAT}}}^{1}_{3}$.
Proof.
We implement Alg.˜3 on ${{{\textsc{AHAT}}}}s$ and show how linearity reduces the computational requirements w.r.t. Thm.˜4.1. We define $\mathcal{V}$} and $\mathcal{T}$} as in §˜4.1. Assuming linearity, there is an edge from $v_{1}$ to $v_{2}$ if and only if $v_{1}$ takes the form ${{\mathrm{A}}},{{i}},{{j}}]$, $v_{2}$ takes the form ${{\mathrm{B}}},{{i}}+1,{{j}}]$ such that ${{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}{{\mathrm{B}}}}}\in{{\mathcal{P}}}$ (or the symmetric case). We will now prove how assuming linearity reduces both the looping and padding requirements.
We first remark that we now have a constant number of outgoing edges for each node. Due to linearity, rules that spawn non-terminals are of the form ${{\mathrm{A}}}\rightarrow{\texttt{w}}{{\mathrm{B}}}$}} or ${{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{\texttt{w}}$}}, and solving an item ${{\mathrm{A}}},{{i}},{{j}}]$ therefore reduces to solving items that aim to derive either ${\texttt{w}}_{{{i}}+1}\ldots{\texttt{w}}_{{j}}$ or ${\texttt{w}}_{{{i}}}\ldots{\texttt{w}}_{{{j}}-1}$. There are finitely many such items given ${{\mathrm{A}}},{{i}},{{j}}]$ as the indices are fixed. Therefore, the procedure can be implemented with ${{\mathcal{O}}}({{n}}^{2})$ padding tokens.
Moreover, because every production rule now necessarily spawns a terminal symbol, the full dependency graph can be constructed via ${\textsc{DG}}({\mathcal{T}})$. If ${{\mathrm{A}}}\rightarrow{\texttt{w}}{{\mathrm{B}}}$}} is a production rule used in the derivation of a string, then ${\texttt{w}},{{i}},{{i}}]\in{\mathcal{T}}$ for some i{{i}}, and ${{{\mathcal{R}}}}([{\texttt{w}},{{i}},{{i}}],[{{\mathrm{B}}},{{i}}+1,{{j}}],[{{\mathrm{A}}},{{i}},{{j}}])={\mathrm{T}}$. Crucially, any production rule applied in the derivation of a string that reduces some item ${{\mathrm{A}}},{{i}},{{j}}]$ to another item ${{\mathrm{B}}},{{i}}+1,{{j}}]$ leads to an edge between their associated items in the initial dependency graph ${\textsc{DG}}({\mathcal{T}})$. Therefore, we can compute the realizability of all items with a single call to Reach on the initial dependency graph ${\textsc{DG}}({\mathcal{T}})$, and $\log({{n}})$ looping layers then suffice to perform Alg.˜3. ∎
5 Experiments
We conduct experiments to elicit the impact of looping when recognizing CFLs, and provide more details on our experimental setup in App.˜C. We train transformer classifiers on CFLs of varying degrees of complexity, none of which require extra padding for recognition:
  • Boolean formula value problem (BFVP): The set of variable-free Boolean formulas that evaluate to $\mathrm{T}$}. We have proven (Cor.˜4.1) that $\log$-depth and no padding suffice to recognize this language with transformers. We consider formulas in both the standard infix notation (e.g., $1\lor 0$ is in infix notation) and postfix notation (e.g., 11 0 $\lor$ is in postfix notation). Parallel algorithms for BFVP typically rely on postfix notation (buss-bfvp; buss-optimalparallel).
  • Palindrome: The language ${\mathbb{L}}={{\{{{\bm{w}}}{{\bm{w}}}^{R}\mid{{\bm{w}}}\in{{{{\Sigma}}^{*}}}\}}}$ for some alphabet $\Sigma$}}. We focus on a binary alphabet. This language is linear unambiguous and non-deterministic. Prior work has shown that fixed-depth transformers with hard attention can recognize this language (hao-etal-2022-formal).
  • Marked Palindrome: This language simplifies Palindrome by extending strings with a marker #\# between $\bm{w}$}} and ${{\bm{w}}}^{R}$, which delimits at which index we reverse the string. In other words, ${\mathbb{L}}={{\{{{\bm{w}}}\#{{\bm{w}}}^{R}\mid{{\bm{w}}}\in{{({{\Sigma}}/{{\{\#\}}})^{*}}}\}}}$. As a result of the delimiter, this language is linear deterministic.
  • Dyck: The language of nested strings of parentheses of $k$ types, which we denote by $\mathrm{D}\!\left({{k}}\right)$}}. We consider $\mathrm{D}\!\left(1\right)$}} and $\mathrm{D}\!\left(2\right)$}}. This language is non-linear and deterministic. Fixed-depth transformers can recognize $\mathrm{D}\!\left({{k}}\right)$}} for any k{{k}} (weiss2021thinkingliketransformers).
These languages vary in complexity, allowing us to test transformers’ ability to learn CFL recognition constructions for languages of different difficulties. In particular, while Palindrome and $\mathrm{D}\!\left(k\right)$}} languages can in principle be recognized by constant-depth transformers, BFVP requires growing depth (i.e., $\log$-depth), assuming ${\textsc{TC}\textsuperscript{0}}\neq{\textsc{NC}\textsuperscript{1}}$. This suggests that the performance of log-depth vs. constant-depth transformers on BFVP is a good measure of whether transformers can utilize the extra expressivity of log-depth when it is required. Our results are presented in Tab.˜2.
Results. We first highlight that for both variants of BFVP, looping leads to slight improvements in in-distribution (1-3%) and generalization (3-4%) accuracy. We treat BFVP as a testbed of our theory as it is a language known to require $\log$-depth. As we see improvements in performance with looping, our results align with the theory. However, the results are mixed for other languages. For Palindrome and $\mathrm{D}\!\left(1\right)$}}, we notice looping does not improve accuracy. This is supported by the fact that these languages already have fixed-size solutions (hao-etal-2022-formal; weiss2021thinkingliketransformers), and therefore extending the model with dynamically-scaling layers may hinder the performance. However, for $\mathrm{D}\!\left(2\right)$}} and Marked Palindrome, we remark that looping seems to improve generalization even though these languages also have constant-depth transformer constructions. This is supported by the fact that these languages are inexpressible in C-RASP (yang2024countingliketransformerscompiling; huang2025a), a language class that matches closely the set of languages transformers should be able to generalize on.
6 Discussion and Conclusion
Transformers parse in parallel.
Our work provides a theoretical framework for understanding how transformers can internally process syntax: The parsing problem can be formulated as a parallel procedure implementable by looped- and padded-transformers (§˜3, §˜4). Interestingly, transformers in practice seem to implement some form of parallel parsing. schulz2025unravelingsyntaxlanguagemodels show that transformers parse by learning sub-grammars—grammars that generate substrings of the original grammar—in parallel.
Most significantly, allen-zhu2025physics show via probing that transformers simulate a dynamic program that manipulates items: They found that transformers encode the necessary items (of the form ${{\mathrm{X}}},{{i}},{{j}}]$) to parse a string, and that they implement memory reads across positions to combine the solutions of items. While they state that such an algorithm can be naively implemented in polynomial time, our constructions leverage transformers’ inherent parallelism to show it can be implemented exponentially faster (i.e., in logarithmic time). Akin to how transformers can implement state tracking via parallel simulation of automata (liu2023transformerslearnshortcutsautomata; li2025how), our work provides a rigorous construction for how looped transformers can implement parallel parsing which complements recent interpretability results.
On dynamically-scaling transformers.
Fixed-size transformers cannot robustly solve the general CFL recognition problem, neither theoretically (merrill-etal-2022-saturated) nor empirically (khalighinejad-etal-2023-approximating; butoi-languageclasses). There is a simple explanation for this phenomenon: general CFL recognition is an NC1-hard problem (venkateswaran-logcfl). While transformers used in practice are fixed-size, our theoretical analysis can guide novel model architectures with improved algorithmic capabilities over any input.
Learnability.
When aiming to predict the exact empirical capabilities of some parametrized model, expressivity results cannot paint the full picture. A function that is expressible may not be easily learnable (hahn-rofin-2024-sensitive). While our work provides a framework for understanding how transformers can express a context-free language, we encourage future work to further investigate which class of grammars are provably learnable (huang2025a), and if looping and padding can enhance learnability (fan2025loopedtransformerslengthgeneralization).
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments
We thank attendees of Formal Languages and Neural Networks (FLaNN) seminar for insightful discussions about this work. Anej Svete is supported by an ETH AI Center Doctoral Fellowship. William Merrill was supported by a Two Sigma PhD fellowship, an NSF Graduate Research Fellowship, and the Allen Institute for Artificial Intelligence.
Appendix A Extended Background
A.1 Transformer Models
We introduce in this section our idealization of the transformer architecture.
A.1.1 Fixed-size Transformers
An L{L}-layer transformer of constant width  [^5] To guarantee the transformer width is constant while the number of layers grows with input length, we recall transformer layers can reset intermediate values in looping layers (merrill-2025littledepthgoeslong). D{{D}} is a mapping ${{\textsf{T}}}\colon{{{{\Sigma}}^{*}}}\to{{\left({{\mathbb{R}}}^{{D}}\right)^{*}}}$:
${{\textsf{T}}}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{\mathrm{\mathcal{L}}}^{({L})}\circ\cdots\circ{\mathrm{\mathcal{L}}}^{(1)}\circ{\texttt{embed}}$ (3)
The input encoding function ${\texttt{embed}}\colon{{{{\Sigma}}^{*}}}\to{{\left({{\mathbb{R}}}^{{D}}\right)^{*}}}$ applies an injective position-wise embedding function to each symbol in the input string $\bm{w}$}}. We use bos and eos symbols, distinct symbols that are placed at the beginning and end of every input string, respectively.
${\mathrm{\mathcal{L}}}^{({\ell})}$ for ${\ell}\in{{\left[{L}\right]}}$ denotes a transformer layer—a mapping ${\mathrm{\mathcal{L}}}^{({\ell})}\colon{{\left({{\mathbb{R}}}^{{D}}\right)^{*}}}\to{{\left({{\mathbb{R}}}^{{D}}\right)^{*}}}$ that updates the symbol representations. The components of a transformer layer are the layer normalization LN, the attention layer ${{{\bm{f}}}_{\texttt{att}}}^{({\ell})}$ and the feedforward network ${\mathbf{F}}^{({\ell})}$. Concretely:
${\mathrm{\mathcal{L}}}^{({\ell})}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{\mathbf{F}}^{({\ell})}\circ{{{\bm{f}}}_{\texttt{att}}}^{({\ell})}\circ{\texttt{LN}}^{({\ell})}$ (4)
We recall layer-normalization maps a vector ${{{\bm{x}}}}\in{{\mathbb{R}}}^{{n}}$ of some dimension n{{n}} to $\frac{{{{\bm{x}}}}^{\prime}}{{{\left\lVert{{{\bm{x}}}}^{\prime}\right\rVert}}}$ where ${{{\bm{x}}}}^{\prime}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{{\bm{x}}}}-\frac{\sum_{{{{x}}}_{{i}}\in{{{\bm{x}}}}}{{{x}}}_{{i}}}{{{n}}}$. We assume multi-pre-norm (merrill2024expressivepowertransformerschain). In standard pre-norm, we apply a layer-normalization to the entire hidden state of each symbol. In multi-pre-norm, we allow each sublayer to take $k$ different projections of its input apply layer-norm to each and concatenate. Crucially, multi-pre-norm allows us to partition the hidden state and normalize disjoint subsets thereof, which we will rely on in our proofs.
${\mathbf{F}}^{({\ell})}\colon{{\left({{\mathbb{R}}}^{{D}}\right)^{*}}}\to{{\left({{\mathbb{R}}}^{{D}}\right)^{*}}}$ is a position-wise function that applies the same feedforward network to every symbol of the sequence. It is parametrized by weight matrices of the form ${{{\bm{W}}}}\in{{\mathbb{R}}}^{{{m}}\times{{D}}}$ and ${{{\bm{U}}}}\in{{\mathbb{R}}}^{{{D}}\times{{m}}}$. A feedforward network ${\mathbf{F}}^{({\ell})}$ can nest functions of the form ${{{\bm{U}}}}{{\mathrm{ReLU}}}({{{\bm{W}}}}{{{\bm{z}}}})$ where ${{{\bm{z}}}}\in{{\mathbb{R}}}^{{D}}$ is an intermediate value.
The attention mechanism is defined by the function ${{{\bm{f}}}_{\texttt{att}}}^{({\ell})}\colon{{\left({{\mathbb{R}}}^{{D}}\right)^{*}}}\to{{\left({{\mathbb{R}}}^{{D}}\right)^{*}}}$. We denote by ${{{\bm{k}}}}_{{{i}}}^{({\ell})}$, ${{{\bm{q}}}}_{{{i}}}^{({\ell})}$, ${{{\bm{v}}}}_{{{i}}}^{({\ell})}$ the key, query and value vectors, respectively, for symbol i{{i}} at layer $\ell$}. ${{{\bm{f}}}_{\texttt{att}}}^{({\ell})}$ is defined as follows:
${{{\bm{f}}}_{\texttt{att}}}^{({\ell})}(({{{x}}}_{1},\cdots,{{{x}}}_{{{T}}}))$ $\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}({{{y}}}_{1},\cdots,{{{y}}}_{{{T}}})$ (5a)
${{{y}}}_{{i}}$ $\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{{x}}}_{{i}}+\sum_{{{i}}^{\prime}\in{{m}}\left({{i}}\right)}{{{s}}}_{{{i}}^{\prime}}{{{\bm{v}}}}_{{{i}}^{\prime}}^{({\ell})}$ (5b)
${{{s}}}$ $={{\texttt{proj}}}({{\{{{{\texttt{score}}}}({{{\bm{k}}}}_{{{i}}^{\prime}}^{({\ell})},{{{\bm{q}}}}_{{{i}}}^{({\ell})})\}}})$ (5c)
${{m}}\left({{i}}\right)$ is a set that defines the masking used by the transformer. For instance, ${{m}}\left({{i}}\right)={{\{{{i}}^{\prime}\mid{{i}}^{\prime}<{{i}}\}}}$ refers to strict causal masking and ${{m}}\left({{i}}\right)={{\left[|{{\bm{w}}}|\right]}}$ refers to no masking. score is a scoring function that maps two vectors of the same size to a scalar. Typically, the dot-product score is used with ${{{\texttt{score}}}}({{{x}}}_{1},{{{x}}}_{2})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\langle{{{x}}}_{1},{{{x}}}_{2}\rangle$.
Throughout layers, the hidden state ${{{y}}}_{{i}}$ of a symbol at position i{{i}} continuously evolves as it cumulatively adds up the outputs of the attention mechanism. We call this cumulative sum ${{{y}}}_{{i}}$ over layers the residual stream at i{{i}}.
proj is a projection function that normalizes the scores into weights for the symbol values. Following previous work, we assume an averaging hard attention transformer (AHAT), which concentrates the attention weights on the symbols that maximize the attention score (merrill-etal-2022-saturated; strobl2023averagehardattentiontransformersconstantdepth). Formally, we have ${{\texttt{proj}}}={{\mathrm{hardmax}}}$:
Definition A.1.
Averaging hard attention is computed with the $\mathrm{hardmax}$}} projection function:
[…]
for ${{d}}\in{{\left[{{D}}\right]}}$, where ${{{\bm{x}}}}\in{{\mathbb{R}}}^{{D}}$ and $m\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}|\operatorname*{{{argmax}}}\left({{{\bm{x}}}}\right)|$ is the cardinality of the argmax set.
Recognition.
A transformer is a vector-valued function. To link this to language recognition, we use the representations computed by a transformer for binary classification of strings. We denote by ${{{\bm{x}}}}_{{{\textsc{eos}}}}^{{L}}$ the hidden state of eos at the end of the forward pass of T. Typically, string recognition is based on ${{{\bm{x}}}}_{{{\textsc{eos}}}}^{{L}}$ as eos is the only symbol that is able to access information about every single symbol throughout all (assuming causal masking). This allows us to define a transformer’s language based on a linear classifier:
${{\mathbb{L}}}\left({{\textsf{T}}}\right)\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{\{{{\bm{w}}}\in{{{{\Sigma}}^{*}}}\mid{{{\bm{\theta}}}}^{\top}{{{\bm{x}}}}_{{{\textsc{eos}}}}^{{L}}>0\}}}.$ (7)
Precision.
Following previous work (merrill2025exactexpressivepowertransformers; merrill2024expressivepowertransformerschain; merrill-logic), we assume $\log$-precision transformers, i.e., we allow the transformer to manipulate values that can be represented with ${{\mathcal{O}}}(\log({{n}}))$ bits for an input of length n{{n}}. It is a minimally extended idealization that enables the transformer to store indices and perform sums over an unbounded number of symbols, two crucial capabilities for our constructions.
A.1.2 Layer-norm hash
We will often use the layer-norm hash building block (merrill2024expressivepowertransformerschain). It is particularly useful for equality checks between values across different symbols, especially with a potentially unbounded number of queries and keys.
Definition A.2 (merrill2024expressivepowertransformerschain).
Given a scalar ${{{z}}}\in\mathbb{R}$, its layer-norm hash is $\phi({{{z}}})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\langle{{{z}}},1,-{{{z}}},-1\rangle/\sqrt{{{{z}}}^{2}+1}$.
Layer-norm hash is scale invariant, and $\phi(q)\cdot\phi(k)=1$ if and only if $q=k$. In other words, the inner product of scalars $q$ and $k$, even if computed at different positions $i$ and $j$, respectively, allows us to check for the equality of $q$ and $k$. Layer-norm hash therefore allows us to perform equality checks over elements of residual streams at different positions.
Appendix B Transformer Constructions Proofs
In our constructions, we leverage padding tokens to associate them with distinct objects. For example, when computing the realizability of items in Alg.˜1 and Alg.˜2 on ${{{\textsc{AHAT}}}}s$, we will associate each item with a padding token. To this extent, we introduce a novel theoretical gadget implementable by ${{{\textsc{AHAT}}}}s$ that enables a padding token at some position i{{i}} to compute the encoding of its associated items from the unique position i{{i}}. We formalize this statement in the following lemma:
Lemma B.1 (Converting a padding token position into a binary representation).
Let T be a ${{\mathcal{O}}}(\mathcal{P}({{n}}))$-padded transformer. Let $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{1}\ldots\mathcal{S}_{m}$ be some set such that its elements can be represented with ${{\mathcal{O}}}(\log(\mathcal{P}({{n}})))$ bits. Then, in a constant number of layers, each padding token can add to their residual stream the encoding of a distinct element of $\mathcal{S}$.
Proof.
Firstly, a padding token at position i{{i}} can add to the residual stream $\phi({{i}})$ with one causally-masked attention layer by uniformly attending over the strict left context and setting as value $\mathbbm{1}[{{i}}=0]$ (merrill2024expressivepowertransformerschain).
Each padding token is distinguished by its unique position. We will rely on this fact to unpack bits of the binary representation of $\phi({{i}})$ to store the encoding of a distinct element of $\mathcal{S}$.
Recall AHATs can compute Euclidean divisions and modulo at some position i{{i}} for integers smaller than i{{i}} in a constant number of layers (merrill-2025littledepthgoeslong). We leverage this theoretical gadget to partition the binary representation of $\phi({{i}})$ into an element of $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{1}\ldots\mathcal{S}_{m}$. As an example, suppose $\mathcal{S}_{1}={{\left[{{n}}\right]}}$, and $s_{1}$ is some index in $\mathcal{S}_{1}$. $s_{1}$ can then be written with $\log({{n}})$ bits. We can extract $s_{1}$ from $\phi({{i}})$ by considering the binary representation of the latter and extracting the first $\log({{n}})$ bits or equivalently, computing $\phi({{i}})$ mod n{{n}}. To add to the residual stream the next element $s_{2}\in\mathcal{S}_{2}$, we can clear out the first $\log({{n}})$ bits of $\phi({{i}})$ by dividing $\phi({{i}})$ by n{{n}}. This example illustrates how we can extract from $\phi({{i}})$ an element of $\mathcal{S}$: we iteratively 1) mask the first $\log(|\mathcal{S}_{{i}}|)$ bits from the least significant bit to extract an element of $\mathcal{S}_{{i}}$ and 2) shift the binary representation of $\phi({{i}})$ towards the least significant bit to then extract the following element in $\mathcal{S}_{{{i}}+1}$. ∎
B.1 General CFL Recognition on Transformers
See 3.1
Proof.
We store padding tokens for each possible item (of the form ${{\mathrm{X}}},{{i}},{{j}}]$ or ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$) and each possible way to decompose that item. There are ${{\mathcal{O}}}({{n}}^{6})$ such tokens: In the worst case, we are solving an item ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ and are guessing an item ${{\mathrm{Z}}},p,q]$ that decomposes that problem. Intuitively, if a padding token aims to solve the item ${{\mathrm{X}}},{{i}},{{j}}]$ and holds as decomposition ${{\mathrm{Y}}},{{k}},{{l}}]$, we attend to the padding tokens which solve ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ and ${{\mathrm{Y}}},{{k}},{{l}}]$. Due to Thm.˜3.3, if ${{{{\mathrm{S}}}}},1,{{n}}]$ is realizable then there exists a padding token with associated item ${{{{\mathrm{S}}}}},1,{{n}}]$ such that it will store 1 (denoting realizability) in its residual stream after ${{\mathcal{O}}}(\log({{n}}))$ steps.
We firstly detail how each padding token can add to their residual stream the encodings of their associated item and subsequent decomposition. A padding token at position i{{i}} can add to their residual stream $\phi({{i}})$ with one causally-masked attention layer by attending to their strict left context (merrill2024expressivepowertransformerschain). We define the set $\mathcal{S}=\mathcal{S}_{1}\times\ldots\mathcal{S}_{m}$ as the set of all possible item / decomposition combinations. For instance, (${{\mathrm{X}}},{{i}},{{j}}],[{{\mathrm{Y}}},{{k}},{{l}}]$) is an element of this set, where we will decompose ${{\mathrm{X}}},{{i}},{{j}}]$ into ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ and ${{\mathrm{Y}}},{{k}},{{l}}]$. $\mathcal{S}_{1}$ could contain a set of non-terminals in $\mathcal{N}$}}, $\mathcal{S}_{2}$ could contain a set of indices in $\left[{{n}}\right]$}}, so on and so forth. Finally, we leverage Lem.˜B.1 to add the encodings of these elements in the residual stream. For each padding token we can therefore store its associated item and decomposition.
We will now detail how to compute the realizability of items associated with these padding tokens. We consider items of the form ${{\mathrm{X}}},{{i}},{{j}}]$, solving items of the form ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ follows the same idea.
Padding tokens allocate space for an element of $\{0,1,\bot\}$}}, which describes whether the associated item is non-realizable (0), realizable (1), or not known yet to be realizable ($\bot$). Padding tokens initially all store $\bot$.
Base case:
Items of the form ${{\mathrm{X}}},{{i}},{{j}}]$ are a base case item if i=j{{i}}={{j}}. A feedforward network can for each padding token associated with some ${{\mathrm{X}}},{{i}},{{j}}]$ check that i=j{{i}}={{j}} by adding ij{{i}}-{{j}} to the residual stream and checking ij=0{{i}}-{{j}}=0 using $\mathrm{ReLU}$}} activations [^6] This equality check only works because ij{{i}}-{{j}} is guaranteed to be an integer (yang2025transformercookbook).. With an attention layer, we can then retrieve and add to the residual stream the encoding of the symbol ${\texttt{w}}_{{i}}$ for a given base case item ${{\mathrm{X}}},{{i}},{{i}}]$ as follows. A symbol representation at position i{{i}} can add to its residual stream $\phi({{i}})$ by uniformly attending with a causally-masked attention layer to all symbol representations in the strict left context (merrill2024expressivepowertransformerschain). A padding token associated with ${{\mathrm{X}}},{{i}},{{i}}]$ also stores $\phi({{i}})$. Therefore, via an equality-check via dot product, padding tokens can attend to relevant symbol representations by setting as value the one-hot encoding of the symbol $\llbracket{\texttt{w}}_{{i}}\rrbracket$}}. Finally, a feedforward network can add to the residual stream 1 if ${{\mathrm{X}}}\rightarrow{\texttt{w}}_{{i}}$}} is a valid rule and otherwise 0: A mapping between two finite sets ${{\mathcal{N}}}\times{{\Sigma}}\to{{\{0,1\}}}$ can be computed by a feedforward network.
Induction step:
Recall a padding token stores 1) an item to solve (for instance, ${{\mathrm{X}}},{{i}},{{j}}]$) and 2) a set of objects that enable us to decompose that item (for instance, ${{\mathrm{Y}}},{{k}},{{l}}]$). Given ${{\mathrm{X}}},{{i}},{{j}}]$, ${{\mathrm{Y}}},{{k}},{{l}}]$, a feedforward network adds the encodings of ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ and ${{\mathrm{Y}}},{{k}},{{l}}]$ to the residual stream. Otherwise, if a padding token is associated with ${{\mathrm{X}}},{{i}},{{j}}]$, ${{\mathrm{X}}}\rightarrow{{\mathrm{Y}}}{{\mathrm{Z}}}$}} and k{{k}}, we add ${{\mathrm{Y}}},{{i}},{{k}}-1]$ and ${{\mathrm{Z}}},{{k}},{{j}}]$ to the residual stream via a feedforward network. In the latter case, a feedforward network can also ensure the rule ${{\mathrm{X}}}\rightarrow{{\mathrm{Y}}}{{\mathrm{Z}}}$}} is in the grammar, and store 0 in the residual stream (denoting non-realizability) if the rule is not in the grammar.
Finally, with one attention layer and a feedforward network, we can attend to all padding tokens that aim to solve the first subproblem (${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$) and copy the integer in the allocated cell for realizability. We also perform the same procedure for the second subproblem to solve.
We compute the realizability of the current item via an extension of standard Boolean logic (Tab.˜3) to handle the case where padding tokens have not yet computed the realizability of their associated item. We do not elicit the standard rules of propositional logic for brevity.
[…]
Table 3: Truth table for a three-valued logic
that handles propositions with unknown truth value.
Crucially, a feedforward network can compute this mapping as it is between two finite sets.
After at most $\log({{n}})$ steps, some padding token aiming to solve an item ${{\mathrm{A}}},{{i}},{{j}}]$ will necessarily store $1$ if and only if ${{\mathrm{A}}},{{i}},{{j}}]$ is realizable: There exists some balanced decomposition represented by two padding tokens that we can attend to and store the realizability of their associated items.
Recognition step:
The eos token can uniformly attend to all padding tokens that encode the item ${{{{\mathrm{S}}}}},1,{{n}}]$ (we can add ${{{{\mathrm{S}}}}},1$ and n{{n}} to the residual stream beforehand) item and ensure one of them holds 1, denoting realizability. ∎
B.2 Unambiguous CFL Recognition on Transformers
See 4.1
Proof.
We will implement rytter-2pda’s parallel pebble game algorithm for evaluating Boolean formulas in ${{\mathcal{O}}}(\log({{n}}))$ steps. We first formalize different objects we associate with a node. Recall every node $v$ in the binary tree induced by $\psi$ is represented by some padding token which stores pointers to its input arguments. For the padding token associated with node $v$, we allocate space for the following objects:
  • value is the result of evaluating the formula associated with $v$.
  • ptr is a pointer to a node in the computation tree. Initially, all padding tokens store a pointer to themselves. Intuitively, if the value of ptr is known, we can compute the value of the formula associated with $v$.
[…]
The parallel pebbling game consists of three steps which are repeated ${{\mathcal{O}}}(\log({{n}}))$ times: activate, square and pebble. We introduce each operation and detail how to perform them on AHATs.
activate: Recall that $v$’s padding token stores pointers to its input arguments $v_{1}$ and $v_{2}$. If the value of $v_{1}$ is known, ptr is set to $v_{2}$ (and vice-versa). $v$’s padding token can attend to $v_{1}$’s and $v_{2}$’s padding tokens via an equality-check and copy $v_{1}$.value and $v_{2}$.value. Suppose that $v_{1}$’s value is known (the symmetric argument with $v_{2}$ is the same). We will detail how to define $v$’s condf depending on $v_{1}$’s value and $v$’s function type. For instance, if $v$’s function type is $\land$ and $v_{1}$ is known to evaluate to $\mathrm{T}$}, we know $v$’s value is exactly ptr.value, and therefore we define the conditional function as condf(x)=x(x)=x $\forall x\in{{\{0,1\}}}$. We detail all the distinct cases in the following table.
[…]
Table 4: Defining $v$’s relation to ptr’s value depending on $v_{1}$.value and $v$’s function type.
Feedforward networks are able to compute conditional functions (yang2025transformercookbook). Therefore, a feedforward network can add to $v$’s residual stream a pointer to ptr, 0 or 1 depending on the cases presented in Tab.˜4.
square: We then compute the one-step closure of Activate. Let $v$.ptr = $v^{\prime}$ and $v^{\prime}$.ptr = $v^{\prime\prime}$. We first update $v$.ptr with $v^{\prime}$.ptr = $v^{\prime\prime}$ by having $v$’s padding token attend to $v^{\prime}$’s padding token and copy $v^{\prime}$.ptr. Furthermore, by copying $v^{\prime}$’s condf via another attention layer, a feedforward network can compose the conditional functions of $v$ and $v^{\prime}$.
pebble: Finally, we evaluate at the current iteration $v$.value ==condf($v$.ptr.value) as follows. If condf is a constant function, a feedforward network simply modifies $v$.value with a constant value. If condf(x)=x({{{x}}})={{{x}}}, we compute condf($v$.ptr.value) via an equality-check to copy $v$.ptr.value.
We refer to rytter-2pda for the original presentation of this algorithm and the proof of the ${{\mathcal{O}}}(\log({{n}}))$ time bound. ∎
See 4.1
Proof sketch.
We assume the input formula is in postfix notation, which is defined recursively as follows:
  • $\mathrm{T}$} and $\mathrm{F}$} are formulas in postfix notation.
  • If $\alpha$ is a formula in postfix notation then $\neg\alpha$ is a formula in postfix notation.
  • If $\alpha$ and $\beta$ are formulas in postfix notation then $\alpha\beta\land$ and $\alpha\beta\lor$ are formulas in postfix notation.
We first detail how a transformer can determine whether an input formula is well-formed and how it can add to the residual streams of tokens representing an operator $\lor$ (or $\land$, $\neg$) the encodings to the arguments of that operator. To this extent, we will write a C-RASP program (yang2024countingliketransformerscompiling) to compute whether an input formula is well-formed and what are the arguments of each operator in the formula. C-RASP defines a syntax for writing programs, and a C-RASP program defines a formal language. It is known that C-RASP is equivalent to temporal counting logic (yang2024countingliketransformerscompiling), which is a lower bound on the expressive power of fixed-depth average-hard attention transformers (barcelo2024logical). It is therefore established that $\text{{C-RASP}}\subseteq{{\textsc{AHAT}}}^{0}_{0}$.
A C-RASP program consists of a finite sequence of C-RASP operations ${\mathcal{P}}_{1},{\mathcal{P}}_{2},\ldots,{\mathcal{P}}_{l}$. To signal acceptance, the last operation ${\mathcal{P}}_{l}$ is evaluated at the last string position. The atomic C-RASP operations are denoted by ${\pi}_{{a}}({{i}})$, where ${\pi}_{{a}}({{i}})={\mathrm{T}}$ if and only if ${\texttt{w}}_{{i}}={{a}}$. The inductive C-RASP operations include the standard Boolean connectives as well as counting abilities (e.g., counting the number of past positions such that a formula is satisfied, comparing integers).
We first write a C-RASP program that determines if an input formula is well-formed. A formula in postfix notation is well-formed if operators never try to consume more operands than are available and the formula ends with all operands being consumed. We can express the aforementioned procedure with the following C-RASP program.
$\text{{depth}}({{i}})$ $=\#{{j}}\leq{{i}}[{\pi}_{{\mathrm{T}}}({{j}})\lor{\pi}_{{\mathrm{F}}}({{j}})]-\#{{j}}\leq{{i}}[{\pi}_{\lor}({{j}})\lor{\pi}_{\land}({{j}})]$ (8a)
$\text{{well-formed}}({{i}})$ $=[\#{{j}}\leq{{i}}[\text{{depth}}({{j}})<1]=0]\land[\text{{depth}}({{i}})=1]$ (8b)
To determine if a formula is well-formed, well-formed is evaluated at the last position of the input string.
[…]
$\text{{binary-op}}({{i}})$ $=[{\pi}_{\land}({{i}})\lor{\pi}_{\lor}({{i}})]$ (9a)
$\text{{unary-op}}({{i}})$ $={\pi}_{\neg}({{i}})$ (9b)
$\text{{depth}}({{i}})$ $=\#{{j}}\leq{{i}}[{\pi}_{{\mathrm{T}}}({{j}})\lor{\pi}_{{\mathrm{F}}}({{j}})]-\#{{j}}\leq{{i}}[{\pi}_{\lor}({{j}})\lor{\pi}_{\land}({{j}})]$ (9c)
$\text{{dindex}}({{i}})$ $=\#{{j}}\leq{{i}}[\text{{depth}}({{j}})=\text{{depth}}({{i}})]$ (9d)
$\text{{previous}}({{k}},{{i}})$ $=[{{k}}=[[\#{{j}}\leq{{i}}{\mathrm{T}}]-1]]$ (9e)
$\text{{argument}}({{k}},{{i}})$ $=[\text{{unary-op}}({{i}})\land\text{{previous}}({{k}},{{i}})]$ (9f)
$\lor[\text{{binary-op}}({{i}})\land[\text{{previous}}({{k}},{{i}})$ (9g)
$\lor[\text{{depth}}({{k}})=\text{{depth}}({{i}})\land[\text{{dindex}}({{k}})=\text{{dindex}}({{i}})-1]]]$ (9h)
Because $\text{{C-RASP}}\subseteq{{\textsc{AHAT}}}_{0}^{0}$, we can check with a fixed-size transformer whether an input Boolean formula is in postfix notation. Moreover, for every input token associated with an operator, we can add to the residual stream encodings of the tokens that are associated with the operands of that operator via the argument binary predicate. We can therefore invoke Lem.˜4.1 to evaluate this Boolean formula with ${{\mathcal{O}}}(\log({{n}}))$-looping and no additional padding. ∎
See 4.1
Proof.
Each item ${{\mathrm{A}}},{{i}},{{j}}]$ is associated with a padding token. Each potential edge between nodes representing items ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$ is associated with a padding token. There are ${{\mathcal{O}}}({{n}}^{3})$ such padding tokens. We leverage Lem.˜B.1 to enable padding tokens to add to their residual stream the encodings of their associated items from $\phi({{i}})$, the layer-norm hash of their position i{{i}}.
Each padding token for nodes allocates space to store an element in $\{0,1,\bot\}$}} to denote that the associated item is either non-realizable (0), realizable (1) or not known yet to be realizable ($\bot$). We will implement Alg.˜3’s algorithm on AHATs to compute whether items are part of the closure ${\mathcal{C}}({{\bm{w}}})$ (i.e, are realizable) or not.
Initial items. A padding token for some node can check whether its associated item is of the form ${{\mathrm{A}}},{{i}},{{i}}]$ via a feedforward network that checks that the indices are the same. For all such padding tokens, another feedforward network adds $1$ to the residual stream if and only if ${{{{\mathrm{A}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}$ to signal the realizability of that item (and otherwise adds 0). We can perform this procedure exactly as in the base case of §˜B.1.
Creating the dependency graph. Padding tokens for edges store items of the form ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$. There are finitely many ${{\mathrm{C}}},{{k}}+1,{{j}}]$ such that ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$ (proportionally many in $|{{\mathcal{N}}}|$), which can be added to the residual stream via a feedforward network. According to Eq.˜1, we set an edge between nodes associated with ${{\mathrm{A}}},{{i}},{{j}}]$ and ${{\mathrm{B}}},{{i}},{{k}}]$ if and only if there is an item ${{\mathrm{C}}},{{k}}+1,{{j}}]$ such that ${{\mathrm{C}}},{{k}}+1,{{j}}]$ is realizable (i.e, the corresponding padding token stores 1 in its residual stream) and ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$. The padding token for the edge associated with ${{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}]$ can check whether any of the items of the form ${{\mathrm{C}}},{{k}}+1,{{j}}]$ are realizable and satisfies ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$ via an equality-check with an attention layer (to check the realizability of the items) and a feedforward-network (to check whether ${{{{\mathrm{A}}}\rightarrow{{\mathrm{B}}}{{\mathrm{C}}}}}\in{{\mathcal{P}}}$). If such an item exists, the padding token associated with ${{\mathrm{A}}},{{i}},{{j}}]$ and ${{\mathrm{B}}},{{i}},{{k}}]$ signals that there is an edge between them in the dependency graph.
Binarization. To efficiently perform reachability queries on a dependency graph, we require it to be binary (rytter-2pda). To this extent, we define the graph transform ${{\mathcal{T}}}\colon{\mathcal{G}}\to{\mathcal{G}}^{\prime}$ which binarizes a given directed graph $\mathcal{G}$} by adding more edges and nodes. Denoting ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ and ${\mathcal{G}}^{\prime}=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{\mathcal{T}}}({\mathcal{G}})$, our graph transform satisfies ${\mathcal{V}}\subset{\mathcal{V}}^{\prime}$ as it simply adds more nodes to the original graph. We now describe $\mathcal{T}$}}.
Given a node with k{{k}} out-neighbors, $\mathcal{T}$}} adds k{{k}}-2 extra nodes in $\mathcal{G}$$^{\prime}$ to create a right-branching binary tree. We denote by $r$ some root node and $v_{1},v_{2},\ldots v_{{k}}$ the out-neighbors of $v$ from $\mathcal{G}$}. $\mathcal{T}$}} introduces k{{k}}-2 extra nodes, $h_{1},h_{2},\ldots,h_{{{k}}-2}$, that are used as follows. The root node $r$ in $\mathcal{G}$$^{\prime}$ has edges to $v_{1}$ and $h_{1}$. For 0<i<k20<{{i}}<{{k}}-2, we instantiate the edges $(h_{{{n}}-2},v_{{{n}}-1})$ and $(h_{{{n}}-2},v_{{{n}}})$. For i=n2{{i}}={{n}}-2, we instantiate the edges $(h_{{{n}}-2},v_{{{n}}-1})$ and $(h_{{{n}}-2},v_{{{n}}})$. The resulting tree is a binary right-branching tree. Crucially, we get the following fact: If there is a path between two nodes $v_{1},v_{2}$ in $\mathcal{G}$}, there is a path between the corresponding nodes $v_{1},v_{2}$ in ${{\mathcal{T}}}({\mathcal{G}})$. Our construction trivially preserves reachability query results. We now show how to perform $\mathcal{T}$}} on transformers.
We assume that the graph $\mathcal{G}$} is encoded in our transformer akin to Lem.˜4.1: Graph nodes are assigned to tokens, and each token stores the encodings of the subsequent tokens that correspond to neighboring nodes. To perform $\mathcal{T}$}}, we assume that k2{{k}}-2 padding tokens are appended to the input for every node with k{{k}} out-neighbors. We will make use of Gorn addresses (gorn) to identify and encode the nodes of the new graph $\mathcal{G}$$^{\prime}$ with bit string addresses. The addresses are defined recursively. The root node is associated with the empty bit string $\varepsilon$. An arbitrary node associated with the bit string $b_{1}b_{2}\ldots b_{h}$ characterizes the Gorn addresses of its two children with $b_{1}b_{2}\ldots b_{h}0$ and $b_{1}b_{2}\ldots b_{h}1$. For instance, Fig.˜1 shows a right-branching tree with the corresponding Gorn addresses.
$\varepsilon$011011110111
Figure 1: Right-branching binary tree with Gorn addresses as node labels.
We require that a token representing some node in this right-branching tree is assigned the correct Gorn address (for instance, $r$ is assigned $\varepsilon$). We can thus simply assume that the tokens associated with these nodes are ordered such that each token can add to its residual stream the correct integer representation of its Gorn address. We rely on Lem.˜B.1 to compute distinct Gorn addresses from token positions. Then, the novel tokens associated with intermediary nodes $h_{1},h_{2},\ldots$ can add to their residual streams the encodings of their descendants in the binary tree as follows. To compute the Gorn address of the first descendant, we shift towards the left the binary representation of the integer by multiplying the scalar representation by 2 via a feedforward network. We obtain the Gorn address of the second descendant by adding 1 to the integer representation of the Gorn address of the first descendant.
Therefore, by performing the graph transform $\mathcal{T}$}} on transformers, we can binarize a given dependency graph in the context of UCFL recognition. Recall that initially, a token associated with ${{\mathrm{A}}},{{i}},{{j}}]$ could attend to all tokens associated with edges of the form $([{{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}])$ via an equality-check. There are linearly many edges $([{{\mathrm{A}}},{{i}},{{j}}],[{{\mathrm{B}}},{{i}},{{k}}])$, and therefore the initial dependency graph has nodes with linearly many out-neighbors. Via the previous transformer construction, we can assume the input has ${{\mathcal{O}}}({{n}}\times{{n}}^{2})={{\mathcal{O}}}({{n}}^{3})$ extra padding tokens (${{\mathcal{O}}}({{n}})$ extra tokens for every item ${{\mathrm{A}}},{{i}},{{j}}]$) and we can therefore binarize a given dependency graph ${\textsc{DG}}(\mathcal{X})$ for some set of marked nodes $\mathcal{X}$. The total amount of padding tokens used in our construction is still ${{\mathcal{O}}}({{n}}^{3})$.
Solving reachability queries. Because the given grammar is unambiguous, for each node $v$, the subgraph induced by nodes reachable from $v$ becomes a tree rooted at $v$ (Fact˜4.1). Reachability queries over binary trees now reduce to evaluating the Boolean formula associated with the binary tree. Leaf nodes associated with realizable items are assigned $\mathrm{T}$}. A non-leaf node has a path to such a leaf if evaluating the induced Boolean expression where non-leaf compute $\lor$ over their children yields $\mathrm{T}$}. We can therefore invoke Lem.˜4.1 to evaluate this Boolean formula.
Recognition step. The eos token can attend to the padding token for node associated with ${{{{\mathrm{S}}}}},1,{{n}}]$ and check whether it is realizable, i.e., store $1$ in its residual stream. ∎
Appendix C Experimental Setup
Data.
We used butoi-languageclasses’s length-constrained sampling algorithm for CFLs to generate datasets. To sample strings from a given grammar $\mathcal{G}$}}, we consider the probabilistic version of $\mathcal{G}$}} which induces a probability distribution over ${\mathbb{L}}({{\mathcal{G}}})$, which enables length-constrained sampling. Importantly, their procedure first samples a desired string length n{{n}}, and then performs sampling over the distribution of all strings of length n{{n}}. Negative strings are either sampled at random from $\Sigma$$^{*}$}} or are perturbations from positive strings by applying random edits to them, the number of which is randomly sampled from a geometric distribution that favors small values (butoi2025trainingneuralnetworksrecognizers).
We therefore follow this procedure to sample positive and negative strings from handpicked context-free grammars except for BFVP. For BFVP, the negative strings were sampled Boolean formulas that evaluate to $\mathrm{F}$}, such that the transformer is trained to correctly evaluate a Boolean formula rather than checking if an input is well-formed. We argue that the ability to process hierarchically nested structures, such as nested subformulas in BFVP, is already captured by the grammar $\mathrm{D}\!\left({{k}}\right)$}}.
The training set consists of 1 million samples with string length at most 40. The test set has 2000 samples with string length at most 80. Testing the model on strings longer than those seen in training enabled the evaluation of its ability to generalize out-of-distribution.
Models and Training Procedure.
We trained causally masked looped transformers with no positional embeddings. We used the PyTorch implementation of a transformer encoder layer with pre-norm. Following our definition of the transformer in §˜2.2, we instantiated our models with an initial block of 2 transformer layers, a looping block (which is repeated $\log({{n}})$ times or once at inference) of 2 transformer layers and a final block of 2 transformer layers. A binary classifier (2 layer feedforward network) was then applied to the final contextual representation of eos. Our transformers have 1.2 million parameter budget. We used the AdamW optimizer (loshchilov2019decoupledweightdecayregularization) and binary cross-entropy loss, considering runs across 5 different seeds. The batch size was set to 64 and the learning rate to 0.0001.